Qs=20r + 100p
Qd=4000 - 100p + 10 i

N Quiz: written Assignment #1 DUX 1 + Intranet | Sandhils PMenu | Beef, Pork Ch YO. Ranch Steakhou & N About Us . Hein BBC ⓑ Dismal River Ob . 0 words D Question 6 1 pts Using the supply and demand equations for wheat above, determine how the equilibrium price and quantity vary with an increase in rainfallr) holding other factors that influence supply and fixed. 1 pts Question 7 Using the supply and demand equations for wheat above, explain in words how equilibrium quantity changes with an increase in income().
Answer #1

Firstly, for equilibrium price and quantity:

Qs= Qd i.e.20r + 100p = 4000 - 100p + 10i
which gives, 200p = 4000 + 10i -20r or p= 20 + 0.05i - 0.1r

Thus dp/dr = -(20/200) = -0.1

and d/dr (Q*) = d/dr (20r + 100 (20 + 0.05i -0.1r)

= d/dr ( 20r + 2000 + 5i -10r), which gives d(Q*)/ dr = 10